# Yixin Liu1 2  Lie Lu2  Jihui Jin2  Lichao Sun1  Andrea Fanelli2

### v2

@@ A.1 Robustness Augmentations
  • White Noise: Adds random electronic noise. Adds Gaussian noise with std 0.001.
@@
  • Crop: Robustness to audio truncation, padding, and in-batch audio mixing. While San Roman et al. (2024) use this to obtain a localization mask, we implement crop as one of the edits to gain cropping-based robustness. Specifically, we first randomly select $k$ starting points, then modify $T/2k$ consecutive samples in one of four ways:
@@ A.2 Dataset Details
We utilize a mixed dataset of $4100$ hours in total for training, which contains 100.59-hour LibriSpeech Panayotov et al. (2015), 98.53-hour AudioSet Gemmeke et al. (2017), 879.29-hour Free Music Archive Defferrard et al. (2016), 9-hour MusicCaps Agostinelli et al. (2023), and 3,016.43-hour VoxPopuli Wang et al. (2021). Each dataset is described as follows:
@@
LibriSpeech. LibriSpeech Panayotov et al. (2015) is an English speech dataset derived from audiobooks in the LibriVox project. The dataset contains approximately 1000 hours of read English speech sampled at 16 kHz, with careful segmentation and alignment. The audio is paired with transcribed text, making it suitable for speech recognition tasks. We used a 100.59-hour subset of the full 1000-hour dataset for training.
@@
AudioSet. AudioSet Gemmeke et al. (2017) is a large-scale dataset containing 2,084,320 human-labeled 10-second sound clips drawn from YouTube videos. It consists of 632 audio event classes organized in a hierarchical ontology, covering a wide range of sounds, including human and animal sounds, musical instruments, genres, and common environmental sounds. The dataset contains approximately 5,790 hours of annotated audio. For our experiments, we used a 98.53-hour randomly sampled subset of the full dataset for training.
@@
Free Music Archive. FMA Defferrard et al. (2016) is a large-scale, open-source dataset of music tracks with clear licensing. We used the “large” subset containing 879.29 hours of audio data with 106,574 30-second tracks sampled at 44.1kHz. The dataset spans multiple genres, including Rock, Electronic, Experimental, Hip-Hop, Folk, and Instrumental, making it suitable for training robust watermarking models across diverse musical styles.
@@
MusicCaps. MusicCaps Agostinelli et al. (2023) is a dataset of 5.5k music-text pairs (9.47 hours total), with high-quality human-written captions describing musical attributes like genre, mood, instruments, and tempo. The dataset was created by having music experts write detailed captions for a subset of 5.5k music clips from AudioSet, with each caption carefully describing the musical content, instrumentation, style, and other sonic characteristics. We used a subset of 9 hours for training.
@@
VoxPopuli. VoxPopuli Wang et al. (2021) is a large-scale multilingual speech corpus collected from European Parliament event recordings between 2009-2020. It contains approximately 400K hours of unlabeled speech data across 23 languages, 1.8K hours of transcribed speech for 16 languages, and 17.3K hours of speech-to-speech interpretation data. For English specifically, it provides 24.1K hours of unlabeled speech data and 543 hours of transcribed speech from 1,313 speakers. The dataset has been widely used for representation learning, semi-supervised learning, and interpretation tasks. We used a 3,016.43-hour subset that covers balanced speech from 23 languages for training.
Appendix B Limitations and Broader Impact
@@
While our method demonstrates strong performance across various standard audio transformations, there are still some limitations and areas for improvement. First, while we achieve good robustness against most common audio edits, the attribution robustness against challenging operations like speed changes and generative editing could be further improved. As shown in Table 6, for speed changes, our method achieves only 4% average attribution accuracy across different user setups, though this still outperforms existing baselines which completely fail. One potential direction for further boosting robustness against speed variations is to incorporate a temporal adjustment layer, similar to the one implemented in Westerfeld (2020). This layer aims to search in a black-box manner for the optimal speed factor to reverse the change. We provide a preliminary exploration of this direction in App. C.3.1.
From a broader impact perspective, audio watermarking technology plays an important role in protecting intellectual property rights and detecting AI-generated content in a proactive manner. However, our experiments on adversarial perturbations (Table 3) show that our model remains vulnerable if the adversary has sufficient access and query opportunities to the watermarking system. This highlights the importance of securing the trained watermarking model and preventing public exposure. Overall, our work contributes to the development of a more robust watermarking solution, and we hope it will inspire further research in this field.
@@ C.1 Subjective Evaluation with MUSHRA
The MUSHRA protocol is a crowdsourced test in which participants rate the quality of various samples on a scale of 0 to 100. The ground truth is provided for reference. We utilized eight randomly sampled speech and music samples from different sources, each lasting 5 seconds. As part of the study, we included a low anchor, which is a lowpass-filtered version at 3.5kHz. Each sample was evaluated by 16 participants, with a post-hoc filtering process to exclude participants who failed to correctly identify more than two hidden reference samples in their scoring, resulting in 12 valid participants at the end. For comparison, the ground truth samples received an average score of around $98$, while the low anchor’s average score was around $25$.
As shown in Figure 5, our method achieved a score of approximately 91, closely matching AudioSeal. Among all the methods, Audiowmark is the one that achieves the highest score at around $94$. WavMark received a slightly lower score of 88, while TimbreWM scored the lowest, around $80$. These results demonstrate that our watermarking approach maintains a perceptual quality that is competitive with SoTA methods.
@@ C.2 Analysis on the Training Dynamics of Models with Different Architectures
AudioSeal’s fully-disjoint architecture in general show slow convergence in both detection and message-bit decoding. After training for 30k steps, it quickly converge to 99% detection accuracy but the learning of message-bit decoding progress slowly, achieving around 70% message bit accuracy at the end of 50k training steps. Moreover, the perceptual quality of AudioSeal is also degraded more than XAttnMark and WavMark.
Compared to these, our blended architecture demonstrates steady and comprehensive learning - both detection and decoding accuracies improve gradually, ultimately achieving near-perfect performance only with about 19k training steps. Notably, with only 2k training steps, XAttnMark can converge to 99% detection accuracy. This great boost in learning efficiency suggests that our partial parameter-sharing strategy and cross-attention mechanism enables more efficient learning of both tasks.
@@ C.3 Robustness to Standard Edits
We present the full results summarizing detection and attribution robustness against standard edits in Tables 5 and 6. For detection, our method achieves consistently high performance across all transformations, with an average accuracy of 99.19% and AUC of 99.93%, significantly outperforming all baselines. Even for challenging transformations like speed changes and neural codec compression (EnCodec) where other methods mostly fail (accuracy around 50-60%), our approach maintains exceptional detection performance with accuracies of 99.5% and 96.5%, respectively.
@@ C.3.1 Simple Speed Reversion Layer with Black-box Grid Search
We explore a post-hoc simple speed reversion layer with a black-box grid search to further boost the robustness against speed changes. The speed reversion layer is added on top of the watermarking model. The main idea is to reverse the speed change by searching for the optimal speed factor with the guidance of the watermarking model. We obverse that on the unwatermarked audio, changing the speed can not adversarially increase the detection accuracy, while on the watermarked audio, doing speed change will significantly affect the watermark message decoding pattern. We observe that the watermarking model predicts more stable message bits for each 1-second chunk when the speed remains unchanged. However, after applying a speed change, the model predicts more diverse message bits with lower confidence for each chunk. To leverage this insight, given the original speed change space $[\gamma_{\min},\gamma_{\max}]$ and the subsequent search space $[\frac{1}{\gamma_{\max}},\frac{1}{\gamma_{\min}}]$, we design the score function $r(\alpha)$ as the average of the watermarking model’s mean detection score over each chunk ${p(\alpha)}$, and the average standard deviation of the predicted message bits ${\bar{s}_{m(\alpha)}}$. Formally, this is expressed as $r(\alpha)=\frac{1}{L}\sum_{i=1}^{L}\left({p(\alpha)}+{\bar{s}_{m(\alpha)}}\right)$, where $L$ is the number of chunks, $\alpha$ is the speed factor, and $m(\alpha)$ is the predicted message bits. Then we conduct a two-round linear grid search to find the optimal speed factor $\alpha$ that maximizes the score function $r(\alpha)$, with the first round step size as $S_{1}$, and the second round step size as $S_{2}$. The speed change parameter $\gamma_{\min}$ and $\gamma_{\max}$ are set to 0.8 and 1.25, respectively. We conduct experiments on both watermarked and unwatermarked audio to investigate whether speed reversion might cause false positives. For the watermarked case, we report the attribution success rate/detection time, and for the unwatermarked case, we report the FPR. As shown by the results in Table 7, with more fine-grained search space, the speed reversion layer can further improve the robustness against speed changes while maintaining the overall detection efficiency. Meanwhile, for unwatermarked audio, applying the speed reversion layer does not falsely increase the probability of being classified as watermarked by the detection model.
Table 7: Comparison of speed reversion layer performance. We report both the success rate and the average detection time (in seconds) for different user pool sizes $N$. The speed reversion layer significantly improves robustness against speed changes while maintaining efficient detection time.
@@ C.4 Robustness to Generative Edits
We use the DDIM-inversion-based editing approach of ZETA (Text-Based Audio Editing) introduced in Manor & Michaeli (2024) to manipulate the audio with text prompts. We briefly summarize the ZETA editing process in the next section. For more details on this editing process, please refer to Manor & Michaeli (2024).
ZETA. Given the original signal $\bm{x}_{0}$ (waveform or latent space), this approach first conduct guided forward process $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\tilde{\mathbf{e}}_{t},\quad t=1,\ldots,T$, where $\bar{\alpha}_{t}$ is the cumulative variance schedule, and $\tilde{\mathbf{e}}_{t}$ is the independent Gaussian noise. Then a sequence of noise vectors that capture the main characteristics of the source audio is extracted using $\mathbf{z}_{t}=\left(\mathbf{x}_{t-1}-\bm{\mu}_{t}\left(\mathbf{x}_{t}\right)\right)/\sigma_{t},\quad t=T,\ldots,1$, where $\bm{\mu}_{t}(\cdot)$ is a denoising neural network trained to predict the mean component of the noise added at time $t$. Note that a source prompt $p_{\text{src}}$ can be also used at the conditioning layer of the denoising neural network to guide the forward diffusion process. During the editing process, given a target prompt $p_{\text{tgt}}$ and a noisy state $\mathbf{x}_{T}$, a guided diffusion backward process is conducted via $\mathbf{x}_{t-1}=\mu_{t}\left(\mathbf{x}_{t}\right)+\sigma_{t}\mathbf{z}_{t},t=T,\ldots,1$, where the noise vector sequence $\mathbf{z}_{t}$ are the ones extracted from the previous forward process. The target prompt $p_{\text{tgt}}$ is embedded and applied on the cross-attention part that affects the editing trajectory.
We present the full results of robustness against generative edits in Table 8, where we evaluate our method against two state-of-the-art audio generation models: Stable Audio Evans et al. (2024) and AudioLDM2 Liu et al. (2024b). Different forward diffusion steps $T$ are used to create varying strengths of audio manipulation, ranging from 10 to 110. The source prompt $p_{\text{src}}$ is set to “A recording of music”, and the target prompt $p_{\text{tgt}}$ is set to “A recording of EDM music with strong rhythm”. We follow the default settings in the original paper for other hyperparameters.
@@
AudioLDM2111Music version is used here 10 0.5938 (0.9375/0.75) 0.5273 0.5313 (0.8125/0.75) 0.4961 0.9375 (0.9375/0.0625) 0.9492
@@
C.5 Generalization Across Standard Editing with Different Strength
We present the generalization of XAttnMark against transformations of different strength in Figure 7. Our method demonstrates remarkable robustness across a wide spectrum of audio transformations with varying intensities. For lossy compression operations, EnCodec shows detection accuracy of 69.5-95.0% and attribution rates of 70.0-100% depending on bitrate. AAC maintains strong detection (97.5%) across all bitrates with attribution ranging from 53.0-98.0%. MP3 compression achieves consistent detection (97.5%) and perfect attribution (100%) across all bitrates.

### v3

@@ A.1 Robustness Augmentations
  • White Noise: Adds Gaussian noise with standard deviation 0.001.
@@
  • Crop: Robustness to audio truncation, padding, and in-batch audio mixing. While San Roman et al. (2024) uses this to obtain a localization mask, we implement crop as one of the edits to gain cropping-based robustness. Specifically, we first randomly select $k$ starting points, then modify $T/2k$ consecutive samples in one of four ways:
@@ A.2 Dataset Details
We utilize a mixed dataset of $4100$ hours in total for training, which contains 100.59-hour LibriSpeech (Panayotov et al., 2015), 98.53-hour AudioSet (Gemmeke et al., 2017), 879.29-hour Free Music Archive (Defferrard et al., 2016), 9-hour MusicCaps (Agostinelli et al., 2023), and 3,016.43-hour VoxPopuli (Wang et al., 2021). Each dataset is described as follows:
@@
LibriSpeech. LibriSpeech (Panayotov et al., 2015) is an English speech dataset derived from audiobooks in the LibriVox project. The dataset contains approximately 1000 hours of read English speech sampled at 16 kHz, with careful segmentation and alignment. The audio is paired with transcribed text, making it suitable for speech recognition tasks. We used a 100.59-hour subset of the full 1000-hour dataset for training.
@@
AudioSet. AudioSet (Gemmeke et al., 2017) is a large-scale dataset containing 2,084,320 human-labeled 10-second sound clips drawn from YouTube videos. It consists of 632 audio event classes organized in a hierarchical ontology, covering a wide range of sounds, including human and animal sounds, musical instruments, genres, and common environmental sounds. The dataset contains approximately 5,790 hours of annotated audio. For our experiments, we used a 98.53-hour randomly sampled subset of the full dataset for training.
@@
Free Music Archive. FMA (Defferrard et al., 2016) is a large-scale, open-source dataset of music tracks with clear licensing. We used the “large” subset containing 879.29 hours of audio data with 106,574 30-second tracks sampled at 44.1kHz. The dataset spans multiple genres, including Rock, Electronic, Experimental, Hip-Hop, Folk, and Instrumental, making it suitable for training robust watermarking models across diverse musical styles.
@@
MusicCaps. MusicCaps (Agostinelli et al., 2023) is a dataset of 5.5k music-text pairs (9.47 hours total), with high-quality human-written captions describing musical attributes like genre, mood, instruments, and tempo. The dataset was created by having music experts write detailed captions for a subset of 5.5k music clips from AudioSet, with each caption carefully describing the musical content, instrumentation, style, and other sonic characteristics. We used a subset of 9 hours for training.
@@
VoxPopuli. VoxPopuli (Wang et al., 2021) is a large-scale multilingual speech corpus collected from European Parliament event recordings between 2009-2020. It contains approximately 400K hours of unlabeled speech data across 23 languages, 1.8K hours of transcribed speech for 16 languages, and 17.3K hours of speech-to-speech interpretation data. For English specifically, it provides 24.1K hours of unlabeled speech data and 543 hours of transcribed speech from 1,313 speakers. The dataset has been widely used for representation learning, semi-supervised learning, and interpretation tasks. We used a 3,016.43-hour subset that covers balanced speech from 23 languages for training.
Appendix B Discussion on Limitations
@@
While our method demonstrates strong performance across various standard audio transformations, several limitations and opportunities for improvement remain. First, although we achieve robust performance against most common audio edits, our attribution accuracy degrades significantly for challenging operations like speed changes and generative editing. Specifically, as shown in Table 7, XAttnMark achieves only 4% average attribution accuracy for speed changes across different user setups, though this still represents an improvement over existing baselines, which fail completely. To enhance robustness against speed variations, one promising direction is to incorporate a temporal adjustment layer similar to Westerfeld (2020), which performs black-box optimization to identify and reverse speed changes. We present preliminary work in this direction in App. C.3.1. Additionally, the current solution’s attribution robustness against AI-based editing remains an area requiring further improvement.
@@ C.1 Subjective Evaluation with MUSHRA
The MUSHRA protocol is a crowdsourced test in which participants rate the quality of various samples on a scale of 0 to 100. The ground truth is provided for reference. We utilized eight randomly sampled speech and music samples from different sources, each lasting 5 seconds. As part of the study, we included a low anchor, which is a low-pass-filtered version at 3.5kHz. Each sample was evaluated by 16 participants, with a post-hoc filtering process to exclude participants who failed to correctly identify more than two hidden reference samples in their scoring, resulting in 12 valid participants at the end. For comparison, the ground truth samples received an average score of around $95$, while the low anchor’s average score was around $25$. As shown in Figure 7, our method achieved a score of approximately 91, closely matching AudioSeal. Among all the methods, AudiowMark is the one that achieves the highest score at around $94$. WavMark received a slightly lower score of 88, while TimbreWM scored the lowest, around $80$. These results demonstrate that our watermarking approach maintains a perceptual quality that is competitive with state-of-the-art methods.
@@ C.2 Analysis on the Training Dynamics of Models with Different Architectures
AudioSeal’s fully-disjoint architecture in general shows slow convergence in both detection and message-bit decoding. After training for 30k steps, it quickly converges to 99% detection accuracy but the learning of message-bit decoding progresses slowly, achieving around 70% message bit accuracy at the end of 50k training steps. Moreover, the perceptual quality of AudioSeal is also degraded more than XAttnMark and WavMark.
Compared to these, our blended architecture demonstrates steady and comprehensive learning - both detection and decoding accuracies improve gradually, ultimately achieving near-perfect performance only with about 19k training steps. Notably, with only 2k training steps, XAttnMark can converge to 99% detection accuracy. This great boost in learning efficiency suggests that our partial parameter-sharing strategy and cross-attention mechanism enable more efficient learning of both tasks.
@@ C.3 Robustness to Standard Edits
We present the full results summarizing detection and attribution robustness against standard edits in Tables 6 and 7. For detection, our method achieves consistently high performance across all transformations, with an average accuracy of 99.19% and AUC of 99.93%, significantly outperforming all baselines. Even for challenging transformations like speed changes and neural codec compression (EnCodec) where other methods mostly fail (accuracy around 50-60%), our approach maintains exceptional detection performance with accuracies of 99.5% and 96.5%, respectively.
@@
To further clarify our attribution evaluation, particularly regarding the metrics used, we also report the False Attribution Rate (FAR) and compare different definitions of attribution accuracy, aligning with methodologies in AudioSeal (San Roman et al., 2024). In our work, attribution accuracy is defined as the fraction of correctly attributed samples among those successfully detected as watermarked. This metric is equivalent to $1-\text{FAR}$ (False Attribution Rate, as reported in AudioSeal). We adopt this definition because it allows for a decoupled analysis of pure attribution performance from the initial detection performance. In AudioSeal, the attribution accuracy is defined as the fraction of correctly attributed samples among all samples, which couples the performance of both detection and attribution. We present the results on both the MusicCaps and VoxPopuli datasets in Table 5, illustrating performance under these specific metrics. The results suggest that in the speech domain, when the user pool size is relatively small, AudioSeal is slightly better than XAttnMark, and in all the other setups, XAttnMark consistently outperforms AudioSeal and WavMark.
Table 5: Attribution results over edits on MusicCaps (our main setup) and VoxPopuli (align with AudioSeal’s setup in San Roman et al. (2024)). We report the false attribution rate (FAR) and two attribution accuracy metrics (Att. and Det.+Att.). The key difference between our definition of attribution accuracy and AudioSeal’s is that we isolate the attribution performance from the detection performance. The best score is in bold, and the runner-up is underlined.
Dataset Metric Method N=10 N=100 N=1000 N=10000
MusicCaps FAR (%) $\downarrow$ WavMark 14.67 14.92 15.18 16.07
AudioSeal 18.77 25.76 36.63 30.53
XAttnMark 6.40 6.13 6.44 6.56
Acc. (Att.) (%) $\uparrow$ WavMark 85.33 85.08 84.82 83.93
AudioSeal 81.23 74.24 63.37 69.47
XAttnMark 93.60 93.87 93.56 93.44
Acc. (Det.+Att.) (%) $\uparrow$ WavMark 79.69 80.56 79.94 79.06
AudioSeal 80.81 73.56 62.69 68.75
XAttnMark 90.50 87.19 90.75 91.81
VoxPopuli FAR (%) $\downarrow$ WavMark 13.11 13.21 13.96 15.00
AudioSeal 3.71 12.86 18.91 16.31
XAttnMark 5.03 6.01 6.25 7.70
Acc. (Att.) (%) $\uparrow$ WavMark 86.89 86.79 86.04 85.00
AudioSeal 96.29 87.14 81.09 83.69
XAttnMark 94.97 93.99 93.75 92.30
Acc. (Det.+Att.) (%) $\uparrow$ WavMark 78.69 78.34 78.06 76.88
AudioSeal 89.50 86.60 79.83 83.45
XAttnMark 92.45 92.69 92.49 91.08
@@ C.3.1 Simple Speed Reversion Layer with Black-box Grid Search
We explore a post-hoc simple speed reversion layer with a black-box grid search to further boost the robustness against speed changes. The speed reversion layer is added on top of the watermarking model. The main idea is to reverse the speed change by searching for the optimal speed factor with the guidance of the watermarking model. We observe that on the unwatermarked audio, changing the speed can not adversarially increase the detection accuracy, while on the watermarked audio, speed changing will significantly affect the watermark message decoding pattern. We observe that the watermarking model predicts more stable message bits for each 1-second chunk when the speed remains unchanged. However, after applying a speed change, the model predicts more diverse message bits with lower confidence for each chunk. To leverage this insight, given the original speed change space $\gamma_{\min},\gamma_{\max}]$ and the subsequent search space $\frac{1}{\gamma_{\max}},\frac{1}{\gamma_{\min}}]$, we design the score function $r(\alpha)$ as the average of the watermarking model’s mean detection score over each chunk $p(\alpha)$}, and the average standard deviation of the predicted message bits $\bar{s}_{m(\alpha)}$}. Formally, this is expressed as $r(\alpha)=\frac{1}{L}\sum_{i=1}^{L}\left(p_{i}(\alpha)+\bar{s}_{m,i}(\alpha)\right)$, where $L$ is the number of chunks, $\alpha$ is the speed factor, $p_{i}(\alpha)$ is the detection score for chunk $i$, and $\bar{s}_{m,i}(\alpha)$ is the average standard deviation of the predicted message bits for chunk $i$. Then we conduct a two-round linear grid search to find the optimal speed factor $\alpha$ that maximizes the score function $r(\alpha)$, with first-round step size $S_{1}$ and second-round step size $S_{2}$. The speed change parameter $\gamma_{\min}$ and $\gamma_{\max}$ are set to 0.8 and 1.25, respectively. We conduct experiments on both watermarked and unwatermarked audio to investigate whether speed reversion might cause false positives. For the watermarked case, we report the attribution success rate/detection time, and for the unwatermarked case, we report the FPR. As shown by the results in Table 8, with a more fine-grained search space, the speed reversion layer can further improve the robustness against speed changes while maintaining the overall detection efficiency. Meanwhile, for unwatermarked audio, applying the speed reversion layer does not falsely increase the probability of being classified as watermarked by the detection model.
Table 8: Comparison of speed reversion layer performance. We report both the attribution success rate and the average detection time (in seconds) for different user pool sizes NN. The speed reversion layer significantly improves robustness against speed changes while maintaining efficient detection time.
@@ C.4 Robustness to Generative Edits
We use the DDIM-inversion-based editing approach of ZETA (Text-Based Audio Editing) introduced in Manor and Michaeli (2024) to manipulate the audio with text prompts. We briefly summarize the ZETA editing process in the next section. For more details on this editing process, please refer to Manor and Michaeli (2024).
ZETA. Given the original signal $\bm{x}_{0}$ (waveform or latent space), this approach first conduct guided forward process $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\tilde{\bm{e}}_{t},\quad t=1,\ldots,T$, where $\bar{\alpha}_{t}$ is the cumulative variance schedule, and $\tilde{\bm{e}}_{t}$ is the independent Gaussian noise. Then a sequence of noise vectors that capture the main characteristics of the source audio is extracted using $\bm{z}_{t}=\left(\bm{x}_{t-1}-\bm{\mu}_{t}\left(\bm{x}_{t}\right)\right)/\sigma_{t},\quad t=T,\ldots,1$, where $\bm{\mu}_{t}(\cdot)$ is a denoising neural network trained to predict the mean component of the noise added at time $t$. Note that a source prompt $p_{\text{src}}$ can also be used at the conditioning layer of the denoising neural network to guide the forward diffusion process. During the editing process, given a target prompt $p_{\text{tgt}}$ and a noisy state $\bm{x}_{T}$, a guided diffusion backward process is conducted via $\bm{x}_{t-1}=\mu_{t}\left(\bm{x}_{t}\right)+\sigma_{t}\bm{z}_{t},t=T,\ldots,1$, where the noise vector sequence $\bm{z}_{t}$ are the ones extracted from the previous forward process. The target prompt $p_{\text{tgt}}$ is embedded and applied on the cross-attention part that affects the editing trajectory.
We present the full results of robustness against generative edits in Table 9, where we evaluate our method against two state-of-the-art audio generation models: Stable Audio (Evans et al., 2024) and AudioLDM2 (Liu et al., 2024a). Different forward diffusion steps $T$ are used to create varying strengths of audio manipulation, ranging from 10 to 110. The source prompt $p_{\text{src}}$ is set to “A recording of music”, and the target prompt $p_{\text{tgt}}$ is set to “A recording of EDM music with strong rhythm”. We follow the default settings in the original paper for other hyperparameters.
@@
AudioLDM2-Music 10 0.5938 (0.9375/0.75) 0.5273 0.5313 (0.8125/0.75) 0.4961 0.9375 (0.9375/0.0625) 0.9492
@@
C.5 Generalization Across Standard Editing with Different Strengths
We present the generalization of XAttnMark against transformations of different strengths in Figure 9. Our method demonstrates remarkable robustness across a wide spectrum of audio transformations with varying intensities. For lossy compression operations, EnCodec shows detection accuracy of 69.5-95.0% and attribution rates of 70.0-100% depending on bitrate. AAC maintains strong detection (97.5%) across all bitrates with attribution ranging from 53.0-98.0%. MP3 compression achieves consistent detection (97.5%) and perfect attribution (100%) across all bitrates.